# Local Spherical Collapsing Box in Athena++: Numerical Implementation and Benchmark Tests

### v1

@@ 3 Numerical Implementation
Here $\lambda^{M}$ and $\lambda^{0}$ denote the maximum and minimum eigenvalues of Roe’s matrix given above, $v_{x,\mathrm{L}}$ and $v_{x,\mathrm{R}}$ are the velocity component normal to the interface in the left and right-states respectively, and $C_{\mathrm{L}}=c_{\mathrm{s},\mathrm{L}}/l$ , $C_{\mathrm{R}}=c_{\mathrm{s},\mathrm{R}}/l$ are the effective sound speed computed from the left- and right- states. For the HLLC solver, the middle contact wave is also included. To compute the fastest wave speed including the middle wave, the rescaled sound speeds $C_{\mathrm{L}}$ and $C_{\mathrm{R}}$ are also used for estimating the pressure in the intermediate states. In the middle wave speed estimation processes, the pressure terms, $p^{*}$, that originate from the momentum equations are replaced with $p^{*}/l^{2}$.
The spatial reconstruction makes use of the left and right eigenmatrice of the Roe’s matrix in the primitive variables, and is modified accordingly. For isothermal hydrodynamics, the eigenmatrice are identical to the original ones presented in Appendix B of (Stone et al., 2008), with the isothermal sound speed replaced by the effective sound speed $C$ in our case, as illustrated previously. For adiabatic hydrodynamics, the eigenmatrice are again in a different form, and are shown in App. A. We implement this modification for both piecewise linear (PLM) and piecewise parabolic (PPM) spatial reconstructions.
Rescaling the sound speed implies the modification of the Courant–Friedrichs–Lewy (CFL) condition. We further consider the effect of the time-varying sound speed on the stability condition, by applying the leapfrog finite difference equation (FDE) to a harmonic oscillator with a time-varying frequency. Our new CFL condition is thus given by
@@
where $C_{\circ}$ is the CFL number  [^2] Time discretization in Athena++ is performed following the method of lines, for which the stability condition depends on both the number of dimensions and the integration method. In a 1D problem, or in multi-dimensional problems with a strong-stability preserving integrator (e.g. second- and third- order Runge-Kutta, Gottlieb et al., 2009), the stability condition is generally $C_{\circ}\leq 1$. However, for 2D or 3D van Leer integrator, the condition is $C_{\circ}\leq 1/2$., and the original physical sound speed has been replaced by the effective sound speeds $c_{\mathrm{s},\mathrm{H}}$ and $c_{\mathrm{s},\mathrm{V}}$. The effect of the time-varying sound speed is reflected in $\mathrm{v}(\beta_{x})$, $\mathrm{v}(\beta_{y})$, and $\mathrm{v}(\beta_{z})$, with $\beta_{x}=\beta_{y}=-U_{0}/c_{\mathrm{s}}$, and $\beta_{z}=-U_{R0}L_{z}/c_{\mathrm{s}}$. Note that both $\beta$ and $\mathrm{v}(\beta)$ are dimensionless. For a collapsing system ($\beta>0$), which is the focus of this work, we have
@@
For a static system, $\beta=0$, we have $\mathrm{v}(\beta\to 0)=1$, which goes back to the original CFL condition, but with rescaled sound speeds. The case of the expanding system ($\beta<0$), as well as the derivation of the new CFL condition are described in App. B.
@@ 4 Numerical tests
In this section, we present benchmark tests of the local box for nonlinear solutions (§4.1) and linear waves (§4.2). Fiducially, we use Roe’s solver with the van Leer integrator (VL2) and piecewise parabolic spatial reconstruction (PPM) for our showcase tests, with isothermal equation of state (EOS). Other parameters of simulation setups, including simulation box sizes and resolutions, are described in each subsection (§4.1, §4.2, and §4.3) for each individual tests. Additionally, we present convergence studies in §4.4, examining both Roe and HLL solvers, considering various time integrators including VL2, second- and third- order Runge-Kutta (RK2, RK3) integrators, with both PPM and piecewise linear (PLM) reconstructions, and both isothermal and adiabatic EOS (with $\gamma=1.4$) tested. CFL number is set to $C_{\circ}=0.4$ for shear flow tests (§4.1), and $C_{\circ}=0.2$ for wave tests (§4.2). We set the code unit with the initial background (unperturbed) gas density $\rho_{0}=1$, and the physical sound speed $c_{s0}=1$. The length unit of the code, $L_{0}=1$, is set by the unit of the local coordinate system defined in Eq. (1) (see also the right panel of Figure 1). We also define the initial values of the length scales in the local model, $R_{0}\equiv\mathcal{R}(0)$, $L_{z0}\equiv L_{z}(0)$, for convenience. For each numerical test, we compare our numerical results with analytical solutions to validate the accuracy and robustness of our implementation. We briefly summarize the analytical solutions corresponding to the tests in each subsections separately. Details of the analytical solutions are presented in LL23.
@@ 4.1 Nonlinear Solutions
Figure 3: Same as Figure 2, but for the test case of elevator flows.
The vertical flow velocity grows as the simulation box collapses, closely matches the analytical expectation.
@@ 4.1.1 Horizontal shear flows
Consider a local model with a spatially homogeneous density $\rho=\rho(t)$ and a purely horizontal velocity field
@@
This is a horizontal shear flow that varies only in the vertical direction. Without loss of generality we can set $A(0)=1$, and the solution to this local model is given with the evolution of the flow amplitude $A$
@@
yielding the following velocity field at time $t$,
@@
In global geometry, this growth of the horizontal flow velocity is a byproduct of the conservation of angular momentum during the spherical collapse.
Figure 4: 3D visualization of the spatial distribution of the diagonal flow test. Arrows mark the direction of the flow, and the total flow velocity magnitude is mapped to colors (see the colorbar). Left panels are the initial setup ($\mathcal{R}=R_{0}$), right panels are at $\mathcal{R}=0.3R_{0}$. Flow direction varies as the collapse proceeds.
We test this solution by setting a local model of uniform collapse, with $R_{0}=10.0L_{0}$, $U_{0}=-1.0c_{\mathrm{s}0}$, $L_{z0}=10.0L_{0}$, $U_{R0}=0.0$. The initial flow structure is set with $v_{x0}=0.025c_{\mathrm{s}0}$, $v_{y0}=0.0125c_{\mathrm{s}0}$, $k_{z}=4\pi$. Simulation box size is set with $L_{x,\mathrm{sim}}=L_{y,\mathrm{sim}}=0.5L_{0},L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 128 cells per unit length. We note that our choice of simulation box size quickly becomes overly large to be physically meaningful, however this is irrelevant for numerically testing the local model. The initial distribution of this velocity field is shown in the upper panel of Figure 2, presenting a 3D visualization of the simulation box. Arrows mark the direction of the velocity vectors, while colors mark the velocity magnitude.
The middle panel of Figure 2 present the simulation result with the peak shear velocity as a function of $\mathcal{R}$, where the peak shear velocity is measured by the maximum value of the horizontally-averaged flow velocity. As expected, the shear flow is amplified as the collapse proceeds, in excellent agreement with the analytical solution. We further define the relative error with L2 norm of each variable $\mathcal{X}$
@@
where $N$ is the total cell number, $i$ denotes each cell, $\mathcal{X}_{i}$ and $\mathcal{X}_{i,\mathrm{exact}}$ are the numerical result and analytical solution for $\mathcal{X}$, respectively. Note that we use the maximum absolute value of the analytical solution over the domain, $\mathcal{X}_{max,\mathrm{exact}}=\max(\left|\mathcal{X}_{i,\mathrm{exact}}\right|)$, to normalize the relative error, in order to avoid meaninglessly large relative errors when the analytical solution is zero or a small number. The time evolution of $\mathrm{err}(v_{x})$ is displayed in the lower panel of Figure 2, showing that our result matches the analytical solution to machine precision.
@@ 4.1.2 Elevator flows
We consider a second class of nonlinear flows in the local model, with a spatially homogeneous density $\rho=\rho(t)$, and a purely vertical velocity field,
@@
This vertical ‘elevator flow’ has a solution with a temporally evolving flow amplitude,
@@
To test the solution of these elevator flows, we set the local model with $R_{0}=10.0L_{0}$, $U_{0}=0.0$, $L_{z0}=10.0L_{0}$, $U_{R0}=-0.1c_{\mathrm{s}0}L_{0}$. The flow structure is initialized with $v_{z0}=0.05c_{\mathrm{s}0}$ and $k_{x}=8\pi$. We choose a 2D simulation box with $L_{x,\mathrm{sim}}=0.5L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 256 cells per unit length. This initial flow structure is displayed in the upper panel of Figure 3.
The middle panel of Figure 3 presents the simulation result of the peak flow velocity as a function of $L_{z}$. The peak flow velocity is obtained by the maximum value of the vertically-averaged flow velocity. The elevator flow grows with time, closely matching the analytical expectation. We further measure the relative error $\mathrm{err}(v_{z})$ as defined in Eq. (70). The lower panel of Figure 3 of the time evolution of $\mathrm{err}(v_{z})$ shows that our result matches the analytical solution to machine precision.
Figure 5: Test case of purely horizontal soundwave with uniform collapse. Upper panel: space-time plot for vertically averaged relative density variation. Dashed lines show the theoretical expectation for phase evolution, in good agreement with the numerical results. Lower panel: relative amplitude of perturbation $\delta\rho/\rho$ as a function of R from the simulation (red solid line). The wave operates under the WKB regime, and is closely consistent with the theoretical expectation from the WKB approximation solution (black dashed line). The wiggles in the simulation results is not present in the WKB solution, but is consistent with the exact solution presented in Eq. (89).
@@ 4.1.3 Diagonal flows
We now consider a class of local models with spatially homogeneous density, $\rho=\rho(t)$, and non-zero horizontal and vertical velocities. The solution for these diagonal flows combines $A=\left(\mathcal{R}/R_{0}\right)^{-2}$ of the horizontal shear flows and $B=\left(L_{z}/R_{0}\right)^{-2}$ of the elevator flows such that
@@
where, without loss of generality, we have aligned the diagonal flow with the $x$-axis. In the global picture, these diagonal flows can be interpreted as a spiral with the pitch angle evolving with time as collapse proceeds, since the horizontal and vertical components of the flow velocity change independently.
We test this solution by setting up a local model with $R_{0}=10.0L_{0}$, $U_{0}=-1.0c_{\mathrm{s}0}$, $L_{z0}=10.0L_{0}$, $U_{R0}=-0.1c_{\mathrm{s}0}/L_{0}$, with initial flow structure $v_{x0}=0.025c_{\mathrm{s}0}$, $v_{z0}=0.05c_{\mathrm{s}0}$, $k_{y}=8\pi$. The simulation box is set with box size $L_{x,\mathrm{sim}}=L_{y,\mathrm{sim}}=0.5L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, and a resolution of 128 cells per unit length. Figure 4 presents the flow structure at the beginning of the simulation ($\mathcal{R}=R_{0}$), as well as later on at $\mathcal{R}=0.3R_{0}$. Variation of flow direction is observed as collapse proceeds. As $U_{0}>U_{R0}L_{z}$, the horizontal flow component grows faster than the vertical component, causing the flow orientation to evolve toward the horizontal direction. The time evolution of flow amplitude closely matches the analytical solution in both horizontal and vertical directions. The results are largely similar to those shown in §4.1.1 and §4.1.2, and are not presented here.
@@ 4.2 Linear Perturbation (Waves)
We now examine the properties of linear perturbations in the local model. For details of the linear wave solutions refer to §4 and Appendix B of LL23, these are briefly summarized here. The density and velocity perturbations $\delta\rho$ and $\delta\mathbf{v}$ of a diagonally propagating sound wave can be written as
@@
where the (complex) wave amplitudes $(X,\Pi)$ fullfill
@@
Here we have introduced the sound wave frequency
@@
Note that, without loss of generality, we align the wave propagating direction with the $x-z$ plane so that $\delta v_{y}=0$.
The behavior of the wave is characterised by two regimes, depending on the value of the ratio between the wave period and the evolution timescale of the background flow, $t_{\mathrm{bg}}\sim\min\left(\mathcal{R}/\dot{\mathcal{R}},L_{z}/\dot{L}_{z}\right)$. For a slowly varying background flow with $\omega t_{\mathrm{bg}}\gg 1$, the wave is well described by a Wentzel–Kramers–Brillouin (WKB) approximation. The WKB approximation for the density and velocity perturbations $\delta\rho$ and $\delta\mathbf{v}$ of a diagonally propagating sound wave is
@@
For a more general test, the exact solution for the sound waves can be obtained for a given collapse profile. Define a length scale of reference $L=L(t)$, and adopt
@@
where $b=L_{z}/\mathcal{R}$ is the aspect ratio of the local reference frame, and the subscription “0” denotes the initial values. This length scale of reference can be equal to the global radius of the cloud, $L=\mathcal{R}$, which can be achieved by adopting $L_{0}=\mathcal{R}_{0}$, and have either $b=b_{0}$ (constant aspect ratio), or $k_{z}=0$ (purely horizontal waves). Consider a collapse profile in the form of ${L}=L_{0}\left(1-t/t_{\mathrm{c}}\right)^{\beta}$, where $t_{\mathrm{c}}$ denotes the collapse timescale, this profile includes two typical types of collapse profiles, namely uniform collapse (${L}=L_{0}+U_{0}t$) where $\beta=1$, and free-fall profile where $\beta=2/3$. For uniform collapse, the exact solution for density and velocity is
@@
Here we have introduced $\Lambda=\frac{1}{2}\sqrt{1-4\left(c_{s}|\mathbf{k}_{0}|t_{c}\right)^{2}}$, and $\left|\mathbf{k}_{0}\right|$ is the initial amplitude of $\mathbf{k}$.
@@
and $J_{\nu}$ and $Y_{\nu}$ correspond to Bessel functions of the first and second kinds respectively, with order $\nu=1/\left(2|1-\beta|\right)$.
Figure 6: Test case of purely horizontal soundwave with non-uniform collapse profile. Upper panel: space-time plot for vertically averaged relative density variation. Lower panel: relative amplitude of perturbation $\delta\rho/\rho$ as a function of R from the simulation (red solid line), consistent with the analytical solution (black dashed line). The vertical dotted line marks the length scale $L_{\mathrm{freezeout}}$ corresponding to freeze-out regime.
Additionally, given the relation $t_{\mathrm{bg}}=L/|\dot{L}|$ ($\beta\neq 1$), the collapsing system can evolve from the WKB regime ($\omega t_{\mathrm{bg}}\gg 1$) to the freeze-out regime ($\omega t_{\mathrm{bg}}\lesssim 1$), where the wave has a longer period than the background flow timescale, and thus show a roughly static pattern. The corresponding length scale at freeze-out can be obtained by setting $\omega t_{\mathrm{bg}}=1$,
@@
For constant aspect ratio $b$, this shear flow is not a physical phenomenon, but rather simply a consequence of the choice of a spatially anisotropic reference frame. However, with a varying aspect ratio during the collapse, the shear flow generated by the sound wave has physical origin, as it cannot be removed by a time-independent rescaling of the anisotropic reference frame.
These aforementioned wave properties in the local model deserve careful testing numerically. Here we test the solutions for horizontal sound wave with both a uniform and a general power law collapse profile separately, as well as the solution for diagonal soundwaves.

### v2

@@ 3 Numerical implementation
Here, $\lambda^{M}$ and $\lambda^{0}$ denote the maximum and minimum eigenvalues of Roe’s matrix given above, $v_{x,\mathrm{L}}$ and $v_{x,\mathrm{R}}$ are the velocity component normal to the interface in the left and right states, respectively, and $C_{\mathrm{L}}=c_{\mathrm{s},\mathrm{L}}/l$ , $C_{\mathrm{R}}=c_{\mathrm{s},\mathrm{R}}/l$ are the effective sound speed computed from the left and right states. For the HLLC solver, the middle contact wave is also included. To compute the fastest wave speed including the middle wave, the rescaled sound speeds, $C_{\mathrm{L}}$ and $C_{\mathrm{R}}$, were also used to estimate the pressure in the intermediate states. In the middle wave speed estimation processes, the pressure terms, $p^{*}$, that originate from the momentum equations were replaced with $p^{*}/l^{2}$.
The spatial reconstruction makes use of the left and right eigenmatrices of the Roe’s matrix in the primitive variables, and was modified accordingly. For isothermal hydrodynamics, the eigenmatrices are identical to the original ones presented in Appendix B of (Stone et al., 2008), with the isothermal sound speed replaced by the effective sound speed, $C$, in our case, as was illustrated previously. For adiabatic hydrodynamics, the eigenmatrices are again in a different form, and are shown in App. A. We implemented this modification for both piecewise linear (PLM) and piecewise parabolic (PPM) spatial reconstructions.
Rescaling the sound speed implies the modification of the Courant–Friedrichs–Lewy (CFL) condition. We further considered the effect of the time-varying sound speed on the stability condition, by applying the leapfrog finite difference equation (FDE) to a harmonic oscillator with a time-varying frequency. Our new CFL condition is thus given by
@@
where $C_{\circ}$ is the CFL number, [^2] Time discretization in Athena++ was performed following the method of lines, for which the stability condition depends on both the number of dimensions and the integration method. In a 1D problem, or in multi-dimensional problems with a strong-stability preserving integrator (e.g. second- and third- order Runge-Kutta, Gottlieb09), the stability condition is generally $C_{\circ}\leq 1$. However, for a 2D or 3D van Leer integrator, the condition is $C_{\circ}\leq 1/2$. and the original physical sound speed has been replaced by the effective sound speeds, $c_{\mathrm{s},\mathrm{H}}$ and $c_{\mathrm{s},\mathrm{V}}$. The effect of the time-varying sound speed is reflected in $\mathrm{v}(\beta_{x})$, $\mathrm{v}(\beta_{y})$, and $\mathrm{v}(\beta_{z})$, with $\beta_{x}=\beta_{y}=-U_{0}/c_{\mathrm{s}}$, and $\beta_{z}=-U_{R0}L_{z}/c_{\mathrm{s}}$. Both $\beta$ and $\mathrm{v}(\beta)$ are dimensionless. For a collapsing system ($\beta>0$), which is the focus of this work, we have
@@
For a static system, $\beta=0$, we have $\mathrm{v}(\beta\to 0)=1$, which goes back to the original CFL condition, but with rescaled sound speeds. The case of the expanding system ($\beta<0$), as well as the derivation of the new CFL condition, are described in App. B.
@@ 4 Numerical tests
In this section, we present benchmark tests of the local box for nonlinear solutions (Sect. 4.1) and linear waves (Sect. 4.2). Fiducially, we used Roe’s solver with the van Leer integrator (VL2) and PPM spatial reconstruction for our showcase tests, with an isothermal EoS. Other parameters of simulation setups, including simulation box sizes and resolutions, are described in each subsection (Sects. 4.1, 4.2, and 4.3) for each individual tests. Additionally, we present convergence studies in Sect. 4.4, examining both Roe and HLL solvers, considering various time integrators including VL2, second- and third-order Runge-Kutta (RK2, RK3) integrators, with both PPM and PLM reconstructions, and both isothermal and adiabatic EoSs (with $\gamma=1.4$) tested. The CFL number was set to $C_{\circ}=0.4$ for shear flow tests (Sect. 4.1) and $C_{\circ}=0.2$ for wave tests (Sect. 4.2). We set the code unit with an initial background (unperturbed) gas density of $\rho_{0}=1$ and a physical sound speed of $c_{s0}=1$. The length unit of the code, $L_{0}=1$, was set by the unit of the local coordinate system defined in Eq. (1) (see also the right panel of Fig. 1). We also defined the initial values of the length scales in the local model, $R_{0}\equiv\mathcal{R}(0)$, $L_{z0}\equiv L_{z}(0)$, for convenience. For each numerical test, we compared our numerical results with analytical solutions to validate the accuracy and robustness of our implementation. We briefly summarize the analytical solutions corresponding to the tests in each subsections separately. Details of the analytical solutions are presented in LL23.
@@ 4.1 Nonlinear solutions
Figure 3: Same as Fig. 2, but for the test case of elevator flows.
The vertical flow velocity grows as the simulation box collapses, closely matching the analytical expectation.
@@ 4.1.1 Horizontal shear flows
Consider a local model with a spatially homogeneous density, $\rho=\rho(t)$, and a purely horizontal velocity field,
@@
This is a horizontal shear flow that varies only in the vertical direction. Without loss of generality, we can set $A(0)=1$, and the solution to this local model is given with the evolution of the flow amplitude, $A$,
@@
yielding the following velocity field at time, $t$,
@@
In global geometry, this growth in the horizontal flow velocity is a byproduct of the conservation of angular momentum during the spherical collapse.
Figure 4: 3D visualization of the spatial distribution of the diagonal flow test. Arrows mark the direction of the flow, and the total flow velocity magnitude is mapped to colors (see the color bar). Left panels are the initial setup ($\mathcal{R}=R_{0}$); right panels are at $\mathcal{R}=0.3R_{0}$. Flow direction varies as the collapse proceeds.
We tested this solution by setting a local model of uniform collapse, with $R_{0}=10.0L_{0}$, $U_{0}=-1.0c_{\mathrm{s}0}$, $L_{z0}=10.0L_{0}$, and $U_{R0}=0.0$. The initial flow structure was set with $v_{x0}=0.025c_{\mathrm{s}0}$, $v_{y0}=0.0125c_{\mathrm{s}0}$, and $k_{z}=4\pi$. The simulation box size was set with $L_{x,\mathrm{sim}}=L_{y,\mathrm{sim}}=0.5L_{0},L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 128 cells per unit length. We note that our choice of simulation box size quickly becomes too large to be physically meaningful; however, this is irrelevant to numerically testing the local model. The initial distribution of this velocity field is shown in the upper panel of Fig. 2, presenting a 3D visualization of the simulation box. Arrows mark the direction of the velocity vectors, while colors mark the velocity magnitude.
The middle panel of Fig. 2 present the simulation result with the peak shear velocity as a function of $\mathcal{R}$, where the peak shear velocity is measured by the maximum value of the horizontally averaged flow velocity. As was expected, the shear flow is amplified as the collapse proceeds, in excellent agreement with the analytical solution. We further defined the relative error with L2 norm of each variable, $\mathcal{X}$,
@@
where $N$ is the total cell number, $i$ denotes each cell, and $\mathcal{X}_{i}$ and $\mathcal{X}_{i,\mathrm{exact}}$ are the numerical result and analytical solution for $\mathcal{X}$, respectively. We used the maximum absolute value of the analytical solution over the domain, $\mathcal{X}_{max,\mathrm{exact}}=\max(\left|\mathcal{X}_{i,\mathrm{exact}}\right|)$, to normalize the relative error, in order to avoid meaninglessly large relative errors when the analytical solution is zero or a small number. The time evolution of $\mathrm{err}(v_{x})$ is displayed in the lower panel of Fig. 2, showing that our result matches the analytical solution to machine precision.
@@ 4.1.2 Elevator flows
We considered a second class of nonlinear flows in the local model, with a spatially homogeneous density, $\rho=\rho(t)$, and a purely vertical velocity field,
@@
This vertical “elevator flow” has a solution with a temporally evolving flow amplitude,
@@
To test the solution of these elevator flows, we set the local model with $R_{0}=10.0L_{0}$, $U_{0}=0.0$, $L_{z0}=10.0L_{0}$, and $U_{R0}=-0.1c_{\mathrm{s}0}L_{0}$. The flow structure was initialized with $v_{z0}=0.05c_{\mathrm{s}0}$ and $k_{x}=8\pi$. We chose a 2D simulation box with $L_{x,\mathrm{sim}}=0.5L_{0}$ and $L_{z,\mathrm{sim}}=1.0L_{0}$, and a resolution of 256 cells per unit length. This initial flow structure is displayed in the upper panel of Fig. 3.
The middle panel of Fig. 3 presents the simulation result of the peak flow velocity as a function of $L_{z}$. The peak flow velocity was obtained by the maximum value of the vertically averaged flow velocity. The elevator flow grows with time, closely matching the analytical expectation. We further measured the relative error, $\mathrm{err}(v_{z})$, defined in Eq. (70). The lower panel of Fig. 3 of the time evolution of $\mathrm{err}(v_{z})$ shows that our result matches the analytical solution to machine precision.
Figure 5: Test case of purely horizontal sound wave with uniform collapse. Upper panel: Space-time plot for vertically averaged relative density variation. Dashed lines show the theoretical expectation for phase evolution, in good agreement with the numerical results. Lower panel: Relative amplitude of perturbation, $\delta\rho/\rho$, as a function of R from the simulation (solid red line). The wave operates under the WKB regime, and is closely consistent with the theoretical expectation from the WKB approximation solution (dashed black line). The wiggles in the simulation results are not present in the WKB solution, but are consistent with the exact solution presented in Eq. (89).
@@ 4.1.3 Diagonal flows
We next considered a class of local models with a spatially homogeneous density of $\rho=\rho(t)$ and nonzero horizontal and vertical velocities. The solution for these diagonal flows combines $A=\left(\mathcal{R}/R_{0}\right)^{-2}$ of the horizontal shear flows and $B=\left(L_{z}/R_{0}\right)^{-2}$ of the elevator flows such that
@@
where, without loss of generality, we have aligned the diagonal flow with the $x$ axis. In the global picture, these diagonal flows can be interpreted as a spiral with the pitch angle evolving over time as the collapse proceeds, since the horizontal and vertical components of the flow velocity change independently.
We tested this solution by setting up a local model with $R_{0}=10.0L_{0}$, $U_{0}=-1.0c_{\mathrm{s}0}$, $L_{z0}=10.0L_{0}$, and $U_{R0}=-0.1c_{\mathrm{s}0}/L_{0}$, with an initial flow structure of $v_{x0}=0.025c_{\mathrm{s}0}$, $v_{z0}=0.05c_{\mathrm{s}0}$, and $k_{y}=8\pi$. The simulation box was set with a box size of $L_{x,\mathrm{sim}}=L_{y,\mathrm{sim}}=0.5L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, and a resolution of 128 cells per unit length. Figure 4 presents the flow structure at the beginning of the simulation ($\mathcal{R}=R_{0}$), as well as later on at $\mathcal{R}=0.3R_{0}$. Variation in the flow direction is observed as the collapse proceeds. As $U_{0}>U_{R0}L_{z}$, the horizontal flow component grows faster than the vertical component, causing the flow orientation to evolve toward the horizontal direction. The time evolution of flow amplitude closely matches the analytical solution in both the horizontal and vertical directions. The results are largely similar to those shown in Sect. 4.1.1 and Sect. 4.1.2, and are not presented here.
@@ 4.2 Linear perturbation (waves)
We examined the properties of linear perturbations in the local model (for details of the linear wave solutions refer to §4 and Appendix B of LL23; these are briefly summarized here). The density and velocity perturbations, $\delta\rho$ and $\delta\mathbf{v}$, of a diagonally propagating sound wave can be written as
@@
where the (complex) wave amplitudes $(X,\Pi)$ fulfill
@@
Here, we have introduced the sound wave frequency
@@
Without loss of generality, we aligned the wave propagating direction with the $x-z$ plane so that $\delta v_{y}=0$.
The behavior of the wave is characterised by two regimes, depending on the value of the ratio between the wave period and the evolution timescale of the background flow, $t_{\mathrm{bg}}\sim\min\left(\mathcal{R}/\dot{\mathcal{R}},L_{z}/\dot{L}_{z}\right)$. For a slowly varying background flow with $\omega t_{\mathrm{bg}}\gg 1$, the wave is described well by a Wentzel–Kramers–Brillouin (WKB) approximation. The WKB approximation for the density and velocity perturbations, $\delta\rho$ and $\delta\mathbf{v}$, of a diagonally propagating sound wave is
@@
For a more general test, the exact solution for the sound waves can be obtained for a given collapse profile. One defines a length scale of reference $L=L(t)$, and adopts
@@
where $b=L_{z}/\mathcal{R}$ is the aspect ratio of the local reference frame, and the subscription “0” denotes the initial values. This length scale of reference can be equal to the global radius of the cloud, $L=\mathcal{R}$, which can be achieved by adopting $L_{0}=\mathcal{R}_{0}$, and which can have either $b=b_{0}$ (constant aspect ratio) or $k_{z}=0$ (purely horizontal waves). If we consider a collapse profile in the form of ${L}=L_{0}\left(1-t/t_{\mathrm{c}}\right)^{\beta}$, where $t_{\mathrm{c}}$ denotes the collapse timescale, this profile includes two typical types of collapse profiles; namely, uniform collapse (${L}=L_{0}+U_{0}t$) where $\beta=1$, and a free-fall profile where $\beta=2/3$. For uniform collapse, the exact solution for density and velocity is
@@
Here, we have introduced $\Lambda=\frac{1}{2}\sqrt{1-4\left(c_{s}|\mathbf{k}_{0}|t_{c}\right)^{2}}$, and $\left|\mathbf{k}_{0}\right|$ is the initial amplitude of $\mathbf{k}$.
@@
and $J_{\nu}$ and $Y_{\nu}$ correspond to Bessel functions of the first and second kinds, respectively, with an order of $\nu=1/\left(2|1-\beta|\right)$.
Figure 6: Test case of purely horizontal sound wave with a nonuniform collapse profile. Upper panel: Space-time plot for vertically averaged relative density variation. Lower panel: Relative amplitude of perturbation, $\delta\rho/\rho$, as a function of R from the simulation (solid red line), consistent with the analytical solution (dashed black line). The vertical dotted line marks the length scale, $L_{\mathrm{freezeout}}$, corresponding to the freeze-out regime.
Additionally, given the relation $t_{\mathrm{bg}}=L/|\dot{L}|$ ($\beta\neq 1$), the collapsing system can evolve from the WKB regime ($\omega t_{\mathrm{bg}}\gg 1$) to the freeze-out regime ($\omega t_{\mathrm{bg}}\lesssim 1$), where the wave has a longer period than the background flow timescale, and can thus show a roughly static pattern. The corresponding length scale at freeze-out can be obtained by setting $\omega t_{\mathrm{bg}}=1$:
@@
For a constant aspect ratio, $b$, this shear flow is not a physical phenomenon, but rather simply a consequence of the choice of a spatially anisotropic reference frame. However, with a varying aspect ratio during the collapse, the shear flow generated by the sound wave has a physical origin, as it cannot be removed by a time-independent rescaling of the anisotropic reference frame.
These aforementioned wave properties in the local model deserve careful testing numerically. Here, we tested the solutions for a horizontal sound wave with both a uniform and a general power law collapse profile separately, as well as the solution for diagonal sound waves.